OpenStudy (anonymous):

find the y-intercepts for the ellipse (x^2 / 4) + (y^2 / 9)

OpenStudy (apoorvk):

if that equation equals 1, then for an ellipse of the form \[\large\frac{x^2} {a^2} + \frac{y^2}{b^2}=1\] the y intercepts are \[+\sqrt{b}\] and \[-\sqrt{b}\]

OpenStudy (anonymous):

so then it would just be the sqrt of 9??

OpenStudy (apoorvk):

both "+" and "-" of sqrt9. since the ellipse cits the Y-axis at two points. And also, if the question asks specifically for the length of the intercept, then fond the distance between the intersection points.

OpenStudy (anonymous):

okay i get that now..